A​ 95% confidence interval for the proportion of U.S. adults who subscribe to some sort of streaming service is​ (0.685, 0.745). The margin of error for this confidence interval​ is:
Answer #1

Solution :

Point estimate = = (Lower confidence interval + Upper confidence interval ) / 2

Point estimate = = (0.685+ 0.745) / 2=0.715

Margin of error = E = Upper confidence interval - = 0.745-0.715=0.03

Margin of error = E =0.03